OpenStudy (anonymous):

Hi! Could someone explain how to solve this? 4^1/3 x 8^1/2 thank you! <3

OpenStudy (anonymous):

4^1/3 = (2^2)^1/3 = 2^2/3 8^1/2 = (2^3)^1/2 = 2^3/2 2^2/3 * 2^3/2 = 2^(2/3 + 3/2) = 2 ^ (4/6 + 9/6) = 2^ (13/6)